Circular No. 8044

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003E IN MCG –4-12-4 M. Papenkova and W. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 17.0) on unfiltered KAIT images taken on Jan. 5.3 and 7.3 UT. SN 2003E is located at α = 4^{h}39^{m}10^{s}.88, δ = −24^{o}10′36′′.5 (equinox 2000.0), which is 5′′.8 west and 9′′.1 north of the center of MCG –4-12-4. A KAIT image from 2002 Nov. 27.3 showed nothing at this position (limiting mag ∼ 18.5).

COMET C/2003 A1 An apparently asteroidal LINEAR object (discovery observation be-

low), posted on the NEO Confirmation Page, has been found to be diffuse by CCD observers elsewhere, including at Haleakala (1.2-m reflector, with K. Lawrence reporting the object as slightly diffuse on NEAT images taken on Jan. 7.3 UT, and again somewhat diffuse on Jan. 8.3), at Kleť (where M. Tichý found a coma diameter of 8′′ on images taken on Jan. 8.7 with the 1.06-m KLENOT reflector), and at Ondřejov (where P. Pravec found a faint, small coma that was “marginally apparent”, on images taken close to the moon on Jan. 8.8 with the 0.65-m f/3.6 reflector).

2003 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 5.07007  1^{h}08^{m}36.^{s}07  −6^{o}18′54′′.8  18.4

MPEC 2003-A56 has astrometry, the following preliminary parabolic orbital elements, and an ephemeris. The object is likely of short period, with the angular orbital elements quite similar to those of D/1783 W1.

T = 2003 Jan. 13.452 TT  ω = 348.^{o}741
Ω = 51.726 }2000.0
q = 2.11485 AU  i = 48.367

2001 QR_{322}

New observations of 2001 QR_{322}, discovered on 2001 Aug. 21 in the course of the Lowell Observatory’s Deep Ecliptic Survey (cf. MPEC 2001-V11), indicate that this object is the first known Neptune ‘Trojan’ object, being a 1:1 Neptune librator that remains more than 20 AU from Neptune over 14000 yr (recent observations and new orbital elements on MPEC 2003-A55). Calculations by E. Chiang, University of California at Berkeley, indicate that its orbit is likely to have remained stable over a billion years.

2003 January 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green